At what temperature in volume at 0°c triple itself, at constant presence

Grade:9

1 Answers

Arun
5 years ago
Dear Rajesh
 
As pressure is constant.
Then
V1/T1 = V2/T2
Let V1 = V
V/273 = 3V/T2
T2 = 3*273 = 819 K
Hence
T2 = 819 K = 819 -273 = 546°C
 
Regards
Arun (askIITians forum expert)